OpenStudy (anonymous):

Two hundred raffle tickets ae sold for $3 each. One prize of $100 is awarded. Raul purchased 1 ticket. Determine his expected value and the fair price for the ticket.

OpenStudy (anonymous):

think of it this way imagine he buys all the tickets, he spends 600 dollars gets back 100 dollars for a net loss of 500 dollars averaged over the total number of tickets the expected value is \[-\frac{500}{200}=-\frac{5}{2}=-2,5\]

OpenStudy (anonymous):

Ok got it....thanks. I would have never guessed it.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

How do I give you a badge?

OpenStudy (anonymous):

you can click "best response" on the right if you like

OpenStudy (anonymous):

There you go....thanks again.

OpenStudy (anonymous):

Still there? I still have to find the fair price for the ticket.

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

okay that was WRONG

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

there are 200 tickets, not 300

OpenStudy (anonymous):

so a fair price would be \(\frac{100}{200}=\frac{1}{2}\) or fifty cents

OpenStudy (anonymous):

Cool, you're the best....

OpenStudy (anonymous):

then you would raise exactly $100 and pay out $100 giving an expected value of zero there is a probability way to do this as well, but it is somewhat more complicated, this is simplest

OpenStudy (anonymous):

I'll take the simple way lol

OpenStudy (anonymous):

yeah, simple is best

OpenStudy (anonymous):

Ok thanks.